calculate the vapour pressure of 10% Equus glucose solution at 30°C if vapour pressure of pure water is 31.8 MM
1
Expert's answer
2020-08-14T07:26:05-0400

"\Delta p=\frac{nb}{na}\times pa"

pa=31.8

10% - 10 g solute in 90 g water

"nb=\frac{m}{M}=\frac{10}{180}=0.056"

"na=\frac{m}{M}=\frac{90}{18}=5"

"\Delta p=\frac{0.056}{5}\times31.8=0.356"